OpenStudy (anonymous):

You are familiar with the following types of factoring: factoring out the Greatest Common Factor (GCF) factoring by grouping factoring trinomials of the form x2 + bx + c and ax2 + bx + c As you know, you need to know the first two types of factoring listed above in order to be successful in factoring trinomials of the form ax2 + bx + c. Part 1: In your own words, explain how a trinomial of the form 2x2 + 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping. Use complete sentences. Part 2: If you were an Algebra 1 instructor and were

OpenStudy (anonymous):

creating a test on factoring trinomials of the form ax2 + bx + c, what do you think would be the easiest way to create a trinomial that can be factored? Provide one unique example.

OpenStudy (anonymous):

\[ax ^{2}+bx+c=0\] find the product ac 1.if ac is positive ,then break b in two smaller parts say d & e such that b=d+e and product de=ac 2. if ac is negative ,then b=f-g,f is greater than b and fg=ac